# Observation of Self-Bound Droplets of Ultracold Dipolar Molecules

### v1

@@ I Introduction
Figure 1: Droplet formation in an ultracold gas of NaCs molecules with non-axially symmetric dipolar interactions. a, A quantum degenerate gas of microwave-dressed NaCs molecules is confined in an anisotropic optical dipole trap. The molecules are simultaneously dressed with a $\pi$ field with linear polarization along the $z$ direction and a $\sigma$ field with elliptical polarization rotating in the $xy$ plane. The frequencies of the dressing fields are close to 3.5 GHz. By tuning the ellipticity $\xi$ of the $\sigma$ field away from pure $\sigma^{+}$ circular polarization, long-range interactions between the molecules take the form of a dipolar interaction for which the axial symmetry is broken. Larger ellipticity |$\absolutevalue{\xi}$ | leads to stronger dipolar interactions. Gravity points along the $-z$ direction. b, Intermolecular interaction potential in the $xy$ plane, calculated for $\xi=-6^{\circ}$. c, Angular dependence of the short-range part (top) and the long-range part (bottom) of the interaction in three-dimensional space for a fixed intermolecular distance ($\xi=-6^{\circ}$). d, Interaction length scales as a function of $\xi$, calculated from the microwave-dressed interaction potential. The solid (dashed) line shows the dipolar length $a_{\mathrm{d2}}$ (s-wave scattering length $a_{\mathrm{s}}$) as defined in the main text. e, Absorption images of a single droplet (left), a weakly dipolar BEC (middle), and a 1D-droplet array (right) after 25 ms time-of-flight expansion. Each image is an average of three individual shots taken along the $z$ direction. Black ellipses (bottom) indicate the trace of the electric field vector of the $\sigma$ field. Scale bar, 30 µm (e).
Here, we report on the observation of self-bound droplets in an ultracold gas of dipolar sodium–cesium molecules. Starting from a weakly interacting molecular BEC, we use microwave dressing fields to ramp up the interaction strength and observe the formation of both single droplets and droplet arrays. By varying the interaction ramp speed over four orders of magnitude, we observe a transition from stable 1D droplet arrays to fluctuating 2D structures, signaling a crossover from near-adiabatic to non-adiabatic state preparation. In the single-droplet phase, we measure an increase of the molecular density by two orders of magnitude compared to the initial BEC. In this regime, the characteristic range of dipolar interactions significantly exceeds the interparticle spacing, indicating the realization of a dipolar many-body system in the strongly interacting regime.
@@ II Strongly Dipolar Interactions
The strength and angular dependence of the dipolar contribution is controlled by the microwave parameters as follows: For a perfectly circularly polarized $\sigma$ field (ellipticity $\xi=0$), the strength of the $\pi$ field can be tuned such that dipolar interactions vanish at long range ($B=0$) as the dipolar interactions induced by the $\sigma$ and $\pi$ fields compensate each other Yuan et al. (2025). Under these conditions, where the intermolecular potential is dominated by the hardcore van der Waals potential, we can efficiently cool molecular samples into a BEC, as described in Ref. Bigagli et al. (2024). While keeping the $\pi$ field fixed at this compensation point, changing the ellipticity of the $\sigma$ field induces a dipolar interaction in which the axial symmetry is broken Syzranov et al. (2014); its anisotropy is illustrated in Fig. 1c and described by Chen et al. (2023)
@@ III Expansion dynamics
For ellipticities between $-3.5(5)^{\circ}$ and $3.5(5)^{\circ}$, the expansion dynamics are consistent with a weakly dipolar BEC Stuhler et al. (2005). At $\xi=0(1)\degree$, where dipolar interactions are minimized, the gas expands into an approximately round cloud, as expected for a non-dipolar, repulsively interacting BEC for our trap parameters at the specific expansion times Ketterle et al. (1999). As $\xi$ deviates from zero, we observe electrostriction: the clouds show a significant elongation along the axis of attractive dipolar interactions.
@@ V Droplet Density
Given the strong contraction of the molecular clouds in the droplet phase, it is of particular interest to measure their density (see Fig. 4). Direct in situ imaging is not feasible, as the droplet size is below our imaging resolution. Instead, we leverage the fast switchability of the microwave fields and utilize the density-dependence of two-body loss ($\propto n$) to infer the molecular density. After preparing the sample at a given ellipticity $\xi$, we abruptly turn off the dressing fields within 10 µs (see Fig. 4a), initiating rapid two-body loss of the unshielded molecules Ospelkaus et al. (2010). The molecular decay dynamics is observed for up to 3 ms; Fig. 4b shows examples for the BEC, droplet, and droplet array phases. In both droplet regimes, the decay is significantly faster than in the weakly dipolar BEC. Assuming that many-body correlations do not significantly modify the loss dynamics, enhanced decay corresponds to a higher molecule density Chomaz et al. (2016); Tanzi et al. (2019). Fitting the observed decay curves using a numerical model based on a local density approximation (see Methods), we obtain the mean density $\langle n\rangle$ for each many-body state, as shown in Fig. 4c.
In the weakly dipolar BEC phase, |$\absolutevalue{\xi}$ | < $3.5\degree$, we measure a mean density of $10^{12}\,\mathrm{cm}^{-3}$, consistent with previous results Bigagli et al. (2024). For $\xi<-3.5\degree$, the density increases sharply, possibly indicating a phase transition between the weakly dipolar BEC and the single droplet phase. For $\xi>3.5\degree$, the density also rises as the system enters the droplet array phase, but the change is more gradual. In the single droplet phase, we measure mean densities up to $\sim 10^{14}\,\mathrm{cm}^{-3}$, an increase by two orders of magnitude compared to the BEC, reaching values that are comparable to densities in ultracold atomic gases. At $\xi=-7.1^{\circ}$, the interparticle spacing is $\sim 4,000\,a_{0}$, approaching the lower bound set by the radius of the repulsive hardcore potential, $\sim 2,000\,a_{0}$. This raises the intriguing question whether the droplet approaches a quantum-liquid or crystalline state Gorshkov et al. (2008); Ciardi et al. (2025). The comparison of the dipolar length, $\sim 13,000\,a_{0}$, with the interparticle spacing yields a ratio of 3.6, indicative of a system in the strongly interacting regime.
@@ VI Conclusions
In conclusion, we have observed self-bound droplets and droplet arrays in a strongly dipolar quantum gas of ultracold molecules. Beyond the expansion dynamics and mean densities discussed here, further investigation of the observed phases will be of great interest. For the strongly interacting single droplet, measurements of vorticity, collective excitations, or shear response could clarify whether the phase has quantum-liquid or crystalline properties Gorshkov et al. (2008); Schmidt et al. (2022); Zhang et al. (2025a); Ciardi et al. (2025). For the droplet arrays, probing phase coherence will elucidate whether the system can show supersolid behavior Tanzi et al. (2019); Böttcher et al. (2019); Chomaz et al. (2019); Norcia et al. (2021); Zhang et al. (2025b). Additional experimental and theoretical studies will be needed to understand the nature of the transitions between the observed phases and the stabilization mechanism of strongly dipolar droplets of microwave-dressed molecules Langen et al. (2025).
@@ VII Acknowledgements
We acknowledge helpful discussions with Blair Blakie, Francesca Ferlaino, Tim Langen, Andrew Millis, Tilman Pfau, Wilhelm Zwerger, and Martin Zwierlein. We thank Lin Su and Tarik Yefsah for critical reading of the manuscript. We are grateful to Aden Lam and Claire Warner for important contributions in the construction of the experimental apparatus. This work was supported by an NSF CAREER Award (Award No. 1848466), an NSF Single Investigator Award (Award No. 2409747), an ONR DURIP Award (Award No. N00014-21-1-2721), an AFOSR Single Investigator Award (Award No. FA9550-25-1-0048), and a grant from the Gordon and Betty Moore Foundation (Award No. GBMF12340). I.S. was supported by the Ernest Kempton Adams Fund. T.K. acknowledges NWO VIDI (Grant ID 10.61686/AKJWK33335). S.W. acknowledges additional support from the Alfred P. Sloan Foundation.

### v2

@@ I Introduction
Figure 1: Droplet formation in an ultracold gas of NaCs molecules with non-axially symmetric dipolar interactions. a, A quantum degenerate gas of microwave-dressed NaCs molecules is confined in an anisotropic optical dipole trap. The molecules are simultaneously dressed with a $\pi$ field with linear polarization along the $z$ direction and a $\sigma$ field with elliptical polarization rotating in the $xy$ plane. The frequencies of the dressing fields are close to 3.5 GHz. By tuning the ellipticity $\xi$ of the $\sigma$ field away from pure $\sigma^{+}$ circular polarization, long-range interactions between the molecules take the form of a dipolar interaction for which the axial symmetry is broken. Larger ellipticity $|\absolutevalue{\xi}$ leads to stronger dipolar interactions. Gravity points along the z-z direction. b, Intermolecular interaction potential in the xyxy plane, calculated for $\xi=-6^{\circ}$. c, Angular dependence of the short-range part (top) and the long-range part (bottom) of the interaction in three-dimensional space for a fixed intermolecular distance ($\xi=-6^{\circ}$). d, Interaction length scales as a function of $\xi$, calculated from the microwave-dressed interaction potential. The solid (dashed) line shows the dipolar length $a_{\mathrm{d2}}$ (s-wave scattering length $a_{\mathrm{s}}$) as defined in the main text. e, Absorption images of a single droplet (left), a weakly dipolar BEC (middle), and a 1D-droplet array (right) after 25 ms time-of-flight expansion. Each image is an average of three individual shots taken along the $z$ direction. Black ellipses (bottom) indicate the trace of the electric field vector of the $\sigma$ field. Scale bar, 30 µm (e).
Here, we report on the observation of self-bound droplets in an ultracold gas of dipolar sodium–cesium (NaCs) molecules. Starting from a weakly interacting molecular BEC, we use microwave dressing fields to ramp up the interaction strength and observe the formation of both single droplets and droplet arrays. By varying the interaction ramp speed over four orders of magnitude, we observe a transition from stable 1D droplet arrays to fluctuating 2D structures, signaling a crossover from near-adiabatic to non-adiabatic state preparation. In the single-droplet phase, we measure an increase of the molecular density by two orders of magnitude compared to the initial BEC. In this regime, the characteristic range of dipolar interactions significantly exceeds the interparticle spacing, indicating the realization of a dipolar many-body system in the strongly interacting regime.
@@ II Strongly Dipolar Interactions
The strength and angular dependence of the dipolar contribution is controlled by the microwave parameters as follows: For a perfectly circularly polarized $\sigma$ field (ellipticity $\xi=0\degree$), the strength of the $\pi$ field can be tuned such that dipolar interactions vanish at long range ($B=0$) as the dipolar interactions induced by the $\sigma$ and $\pi$ fields compensate each other Yuan et al. (2025). Under these conditions, where the intermolecular potential is dominated by the hardcore van der Waals potential, we can efficiently cool molecular samples into a BEC, as described in Ref. Bigagli et al. (2024). While keeping the $\pi$ field fixed at this compensation point, changing the ellipticity of the $\sigma$ field induces a dipolar interaction in which the axial symmetry is broken Syzranov et al. (2014); its anisotropy is illustrated in Fig. 1c and described by Chen et al. (2023)
@@ III Expansion Dynamics
For ellipticities between $3.5(5)^{\circ}$ and $3.5(5)^{\circ}$, the expansion dynamics are consistent with a weakly dipolar BEC Stuhler et al. (2005). At $\xi=0(1)\degree$, where dipolar interactions are minimized, the gas expands into an approximately round cloud, as expected for a non-dipolar, repulsively interacting BEC for our trap parameters at the specific expansion times Ketterle et al. (1999). As $\xi$ deviates from zero, we observe electrostriction: the clouds show a significant elongation along the axis of attractive dipolar interactions.
@@ V Droplet Density
Given the strong contraction of the molecular clouds in the droplet phase, it is of particular interest to measure their density (see Fig. 4). Direct in situ imaging is not feasible, as the droplet size is below our imaging resolution. Instead, we leverage the fast switchability of the microwave fields and utilize the density-dependence of two-body loss ($\propto n$) to infer the molecular density. After preparing the sample at a given ellipticity $\xi$, we turn off the dressing fields within 10 µs (see Fig. 4a), initiating rapid two-body loss of the unshielded molecules Ospelkaus et al. (2010). The molecular decay dynamics is observed for up to 3 ms; Fig. 4b shows examples for the BEC, droplet, and droplet array phases. In both droplet regimes, the decay is significantly faster than in the weakly dipolar BEC. Assuming that many-body correlations do not significantly modify the loss dynamics, enhanced decay corresponds to a higher molecule density Chomaz et al. (2016); Tanzi et al. (2019). Fitting the observed decay curves using a numerical model based on a local density approximation (see Methods), we obtain the mean density $\langle n\rangle$ for each many-body state, as shown in Fig. 4c.
In the weakly dipolar BEC phase, $|\absolutevalue{\xi}$ < $3.5\degree$, we measure a mean density of $10^{12}\,\mathrm{cm}^{-3}$, consistent with previous results Bigagli et al. (2024). For $\xi<-3.5\degree$, the density increases sharply, possibly indicating a phase transition between the weakly dipolar BEC and the single droplet phase. For $\xi>3.5\degree$, the density also rises as the system enters the droplet array phase, but the change is more gradual. In the single droplet phase, we measure mean densities of up to $\sim 10^{14}\,\mathrm{cm}^{-3}$, an increase by two orders of magnitude compared to the BEC, reaching values that are comparable to densities in ultracold atomic gases. At $\xi=-7.1(5)^{\circ}$, the measured mean interparticle spacing is $3,600(900)\,a_{0}$, very close to the lower bound set by the radius of the repulsive hardcore potential, $\sim 2,000\,a_{0}$. This raises the intriguing question whether the droplet approaches a quantum-liquid or crystalline state Gorshkov et al. (2008); Ciardi et al. (2025). The comparison of the dipolar length, $13,000(900)\,a_{0}$, with the interparticle spacing yields a ratio of 4(1), indicative of a system in the strongly interacting regime.
@@ VI Conclusions
In conclusion, we have observed self-bound droplets and droplet arrays in a strongly dipolar quantum gas of ultracold molecules. Beyond the expansion dynamics and mean densities discussed here, further investigation of the observed phases will be of great interest. For the strongly interacting single droplet, measurements of vorticity, collective excitations, or shear response could clarify whether the phase has quantum-liquid or crystalline properties Gorshkov et al. (2008); Cinti and Boninsegni (2017); Schmidt et al. (2022); Zhang et al. (2025a); Ciardi et al. (2025). For the droplet arrays, probing phase coherence will elucidate whether the system can show supersolid behavior Tanzi et al. (2019); Böttcher et al. (2019); Chomaz et al. (2019); Norcia et al. (2021); Zhang et al. (2025b). Additional experimental and theoretical studies will be needed to understand the nature of the transitions between the observed phases and the stabilization mechanism of strongly dipolar droplets of microwave-dressed molecules Langen et al. (2025).
@@ VII Acknowledgements
We acknowledge helpful discussions with Gordon Baym, Blair Blakie, Francesca Ferlaino, Tin-Lun Ho, Tim Langen, Andrew Millis, Tilman Pfau, Dam Son, Frank Wilczek, Wilhelm Zwerger, and Martin Zwierlein. We thank Lin Su and Tarik Yefsah for critical reading of the manuscript. We are grateful to Aden Lam and Claire Warner for important contributions in the construction of the experimental apparatus. This work was supported by an NSF CAREER Award (Award No. 1848466), an NSF Single Investigator Award (Award No. 2409747), an ONR DURIP Award (Award No. N00014-21-1-2721), an AFOSR Single Investigator Award (Award No. FA9550-25-1-0048), and a grant from the Gordon and Betty Moore Foundation (Award No. GBMF12340). I.S. was supported by the Ernest Kempton Adams Fund. T.K. acknowledges NWO VIDI (Grant ID 10.61686/AKJWK33335). S.W. acknowledges additional support from the Alfred P. Sloan Foundation.